Service Observing for DTime UT Date 07062016

Nightlog for 20160706 UT. D-Time plus UAO testing.

On site: David.

Tucson: Jenny, Michelle, Andrew, Shane, Greg.

Local time (MST) used throughout.

19:00 Running MODS1 calibrations, checking LUCI 1 pupil positions.

19:36 Sunset.

20:02 luci1.20160706.0005 is a luci1 pupil image. Only slightly off, so we will use LUCI1 for science.

20:30 1\x94 on the DIMM.

20:33 12 degree twilight

20:35 Starting with LUCI1, Dual AGN programme. Slewing to telluric HD122945. First exposure will be luci1.20160706.0006.

20:41 0\x94.93 on DIMM, 0\x94.6 on guider.

20:42 Starting acquisition

20:50 Attempted to close Aladdin, but all of the LUCI software closed! Starting up the GUIs again.

20:58 Observations were designed for LUCI2, but we are running on LUCI1 due to the DX adsec issue. The 1\x94.5 / 2\x94 slit is being used, but the acquisition positions are quite different on LUCI1. Acquiring on 1082, 610.

21:03 Starting observation. 0009, but it\x92s overexposed. Taking the telluric manually.

21:08 +5\x94 offset to take us off the persistence from the overexposed spectra, taking exp time from 5s to 2.5s.

21:09 Offsetting by -20 and taking the second shot. Good images were 0011 and 0012.

21:10 Slewing to NGC4922.

21:11 18 degree twilight.

21:14 Alarm (at the telescope)! One of the chillers stopped working.

21:16 0\x94.9 on DIMM, 0\x94.76 on guider.

21:17 Starting acquisition. Files 0014 - 0016.

21:25 Starting science. Files 0017 - 0026. Spectra look good.

21:48 Greg arrives. Seeing has been good and stable, we are going to attempt a LUCI1 AO interrupt.

21:52 LUCI1 science observation ends. Setting things up for the interrupt. Loading the modified monocular Stone M script,

22:02 Slewing to Stone M field.

22:04 AO problem, stage timed out. Preset failed.

22:07 Resending preset. No problems this time.

22:09 RunAO command failed. Possibly a pointing issue\x85 Asking David to check.

22:11 Pointing was off by 8\x94 in CA.

22:13 Back on target, starting AO. Guidestar is a double.

22:17 Seeing may be degrading\x85 AO is running. First exposure is 0027.

22:19 Lost AO on first offset. AO star appears to be too faint, possibly due to the degrading seeing.

22:22 Restarting script, re-sending preset.

22:29 Jump in image 0028, presumably due to the double guidestar. GCS is struggling. RunAO failed.

22:35 Restarting script and skipping the second step which we have consistently failed on.

22:41 0029 also shows doubled stars\x85

22:43 Even skipping the troublesome second step we failed to lock the AO on the third step.

22:45 Restarting script, again.

22:51 Skipping to fourth step, having applied ACE preset, and trying again.

22:53 Failed again. AO guidestar is not appearing in the field.

22:55 Giving up on AO interrupt. We plan to stick with LUCI1 science rather than pay for a reconfiguration to MODS and back to LUCI again for UAO tests at 00:30.

22:57 Slewing to SDSSJ135602+182218, LUCI1 Dual AGN programme.

23:00 Pointing is off again, David is correcting.

23:01 LUCI1 is complaining, camera did not move from N30 to N3.75. Resending it from the LUCI1 instrument panel. That cleared the issue.

23:06 Guider 0\x94.85. Starting acquisition, image #0031. There are two targets to put in the slit, but the alignment of the slit if a little off.

23:25 Starting science.

23:42 Cutting this short as we are approaching the elevation limit for the AdSec.

23:42 Too slow! Shell RIP.

23:46 AdSec recovered. LUCI observation execution GUI just went blank. \x91Error in commit Manager\x92. Refreshed the script and tried again with the same result. We could not immediately quite the GUI, it hung for several beats before closing.

23:48 Can not restart the LUCI GUI in the standard way. Restarting the scheduler.

23:52 Tried to slew again with LUCI, no sign of the preset on the TCS side.

23:53 Telescope is now moving. Preset failed.

00:06 Moving to a telluric high enough to be observed.

00:09 Preset failed again, pointing is too far off. David is correcting.

00:11 Trying again. No problems this time.

00:19 Had to close and re-open the LUCI1 readout GUI as it was unresponsive. The script had crossed filters for the through slit image and we had to manually undo this.

00:24 Guider reports seeing of 1\x94.5, 1\x94.4 on the DIMM.

00:27 Handing over to UAO testing.

03:00 UAO work continues. Seeing has degraded significantly and is now worse than 2\x94 and tending towards 3\x94. None the less it seems progress is being made.

03:30 It\x92s too late to do any more science, and UAO tests are ongoing. MODS1 has been put to sleep, and LUCI1 is in a safe state. Astronomy crew are packing up and calling it a night.

03:38 18 degree twilight.

04:15 12 degree twilight.

05:13 Sunrise.

-- %USERSIG{AndrewCardwell - 2016-07-06}%

Comments


 
Topic revision: r1 - 06 Jul 2016, AndrewCardwell
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback